OpenStudy (anonymous):

solve the following equation: a-3/3(a-2)-4/11=0

OpenStudy (anonymous):

(a-3)/(3(a-2))=4/11 cross multiplication 11(a-3)=12(a-2) distribute 11a-33=12a-24 subtract 11a -33=a-24 add 24 -9=a

OpenStudy (anonymous):

Check (-9-3)/3(-9-2)-4/11=0 -12/(-33)-4*3/11*3=0 12/33-12/33=0 tr